OpenStudy (anonymous):

sqr. of 5y - 3x = x - 6y get the dy/dx of this please.

OpenStudy (anonymous):

1/2(5y-3x)^-1/2(5y'-3)=1-6y'

OpenStudy (anonymous):

Now we have to isolate y', usually on left side

OpenStudy (anonymous):

show me how to isolate y" please

OpenStudy (anonymous):

y'

OpenStudy (anonymous):

my answer is \[y' =2 \sqrt{5y-3x} -5y +3 / 6\]